OpenStudy (anonymous):

4a/9=100/45 solve for a

OpenStudy (anonymous):

First, since the right side is being divided by 45, do the opposite and multiply by 45, and what you do to one side you must do to the other, you now you have 45*4a/9=100

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

45*4a= 180a so now your equation is 180a/9=100. Now do the same thing with the 9 and multiply both sides by 9. You will now have 180a=900

OpenStudy (anonymous):